Thorpe Mfg., Inc., is currently operating at only 81 percent of fixed asset capacity. Current sales are $740,000. How fast can sales grow before any new fixed assets are needed? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Current sales at 81% of fixed asset capacity=$740,000

Sales at 100% capacity= 740,000/.81=913,580.24

Maximum Increase in sales without buying extra fixed assets=913,580.24-740,000=$173,580.24

Percentage increase in sales without needing new fixed assets=173,580.24/740,000= .2346 or 23.46%